### v1

# A Systematic Review of NLP for Dementia- Tasks, Datasets and Opportunities

Abstract
The close link between cognitive decline and language has fostered long-standing collaboration between the NLP and medical communities in dementia research. To examine this, we reviewed over 200 papers applying NLP to dementia-related efforts, drawing from medical, technological, and NLP-focused literature. We identify key research areas, including dementia detection, linguistic biomarker extraction, caregiver support, and patient assistance, showing that half of all papers focus solely on dementia detection using clinical data. However, many directions remain unexplored—artificially degraded language models, synthetic data, digital twins, and more. We highlight gaps and opportunities around trust, scientific rigor, applicability, and cross-community collaboration, and showcase the diverse datasets encountered throughout our review—recorded, written, structured, spontaneous, synthetic, clinical, social media-based, and more. This review aims to inspire more creative approaches to dementia research within the medical and NLP communities.
1 Introduction
Dementia is a broad term for a decline in cognitive function caused by various underlying pathologies. It is a progressive, irreversible condition that worsens over time, with no known treatment or cure. The global impact of the disease is staggering. According to the World Health Organization (WHO),1 dementia is currently the seventh leading cause of death globally. As of 2023, approximately 55 million people worldwide are living with dementia, and this number is expected to nearly double every 20 years Prince et al. (2015). Global dementia costs are estimated in the trillions of US dollars, with approximately half of these costs attributed to care provided by informal carers (e.g., family members and close friends), as hospitals and care facilities are overcrowded, and healthcare professionals require specialized training to diagnose patients and address their complex needs.
There are multiple types of dementia, including Alzheimer’s (accounting for 60%-70% of cases111https://www.who.int/news-room/fact-sheets/detail/dementia), vascular dementia, Lewy-body dementia, and others. Although these differ in pathology—some manifesting as abnormal protein deposits and others as reduced blood flow to the brain—they often share common symptoms such as memory loss, confusion, behavioral changes, and language deterioration McKhann et al. (2011). In this review, we use the general term ’dementia’ to encompass all forms and levels of severity of the illness.
Dementia is currently diagnosed through a battery of pathological tests (e.g., brain imaging, blood tests) and cognitive assessments, such as Naming and Verbal Fluency tests, which assess the ability to identify objects or generate words within specific categories Kaplan et al. (2001), and Picture Description tests, evaluating language and narrative skills through prompted descriptions of complex images Mueller et al. (2018). These cognitive assessments are typically conducted in face-to-face interviews with trained medical professionals, who look for linguistic markers of cognitive deterioration, such as repetitive speech, empty speech lacking meaningful content, disorganized narratives, word-finding difficulties, or a tendency to use indirect language and excessive descriptions (e.g., ’the thing you write with’ instead of ’pen’).
Given the prominent role of language in our understanding and diagnosis of the disease, it is no surprise that there is substantial use of NLP methodologies in dementia research within both the NLP and medical communities. Cognitive assessments and similar diagnostics often result in recorded and transcribed data, offering a wealth of textual content that is well-suited for NLP analysis. This, combined with relatively new sources of data (such as social media posts or conversations with LLMs), enhance the potential of NLP to advance dementia research in various directions.
Previous literature reviews on NLP for dementia focus on detection methods (Clarke et al., 2020; Saleem et al., 2022; Qi et al., 2023), with some specifically covering deep learning approaches to detection (Shi et al., 2023; Javeed et al., 2023). Other reviews extend beyond detection, addressing tasks like linguistic biomarker extraction (Gagliardi, 2024) or exploring the role of technology in enhancing patients’ life (D’Onofrio et al., 2017; Shahapure et al., 2022; Saragih et al., 2023; Peres and Campos, 2024), though these are not fully focused on NLP methods. A few reviews cover dementia-related data (De la Fuente Garcia et al., 2020; Yang et al., 2022) or, even more specifically, data from cognitive assessments like picture description tasks (Mueller et al., 2018).
Our review stands out by covering the full spectrum of dementia-related NLP efforts, rather than focusing on specific aspects like data, tasks (e.g., detection), technologies (e.g., deep learning), data types (e.g., clinical), or communities (NLP vs. medical). Moreover, we created our review with NLP readers in mind, unlike prior reviews, which were primarily published in medical literature and aimed at clinicians (see Figure 3). Additionally, we review an extensive 205 papers, covering four scientific communities, categorized by publication venue: the medical community (e.g., Journal of Alzheimer’s Disease); the NLP community (e.g., ACL); the Speech community (e.g., InterSPEECH); and a broader technological community that is not necessarily NLP- or Speech-specific (e.g., Frontiers in Computer Science). In Section 2, we detail the construction of our cohort.
In Section 3 we describe the main research verticals we encountered across our reviewed works, including linguistic biomarker extraction, caregiver support, patient assistance and, most prominently, dementia detection, comprising a striking 57% of the reviewed works focusing on this task alone. For each of these task families, we review motivations, current approaches, and potential future directions.
Our extensive review reveals a broader perspective on the datasets used for these tasks. We find that the vast majority of reviewed works—across all four task families—rely on a few well-known datasets. However, throughout our review we encountered many unique datasets of varying sizes, types, annotations, and purposes. Given the data-driven nature of NLP, we believe highlighting these datasets to the community could significantly advance research and open new avenues for practitioners. We discuss data gaps and opportunities in Section 4, and present a comprehensive table of the 17 datasets we encountered in Table 2.
Furthermore, in Section 4 we offer in-depth insights on the field’s scientific rigor and NLP’s potential to significantly shape medical research, provided trust between the two communities is established. We explore exciting research directions like personalized approaches, artificially degraded language models, and more.
Our novel work primarily aims to inspire researchers from both fields to explore the vast potential of NLP in dementia research. We hope this paper provides a fresh perspective on the domain, emphasizing that the opportunities extend far beyond dementia detection. Whether developing clinical applications, analyzing disease progression, or alleviating caregiver burden, this review serves as a valuable resource for those looking to contribute to the fight against dementia.
2 Methodology
We searched for works applying NLP to dementia-related tasks across six data sources-ACL Anthology, PubMed, DBLP, IEEE Xplore, Springer, and Wiley. For a detailed description of our search queries along with our inclusion and exclusion criteria see Appendix A. In short, our focus was on peer-reviewed papers explicitly mentioning both dementia-related and NLP-focused terms in their titles or abstracts, and that are fully available in English. Many NLP studies on dementia use transcribed speech datasets with both audio and text; we exclude those focusing solely on voice signals without linguistic analysis.
Our screening resulted in 205 relevant papers, which we manually annotated to document their main contributions, datasets used, applied technologies (where applicable), and whether statistical significance was reported (see Appendix A).
We identified four distinct research verticals differing by task and motivation- dementia detection, linguistic biomarker extraction, caregiver support and patient assistance (see Section 3). Two additional, more general categories are literature reviews and dataset introduction papers, which we describe throughout our work. The number of works on NLP for dementia has been steadily growing (Figure 1), reflecting significant interest in the field. While we see a rise in works on patient assistance, caregiver support, and literature reviews, there is still an overwhelming focus on dementia detection compared to other tasks (Figure 2).
In the next section we dive into the four main task families we identify, discussing motivations, trends, and gaps. Naturally, the imbalance between the four tasks (particularly the fact that there are over 110 dementia detection papers while other task families have around 10-40 papers each) affects the depth of our analysis of specific dementia detection studies. Appendix B offers further details on the dementia detection works we reviewed.
3 Task Families
Dementia Detection
Can an algorithm accurately predict dementia from a provided text? Remarkably, 57% of the papers we reviewed address this exact question. This focus stems from two key factors. First, a tangible impact: NLP can significantly improve the diagnostic process patients undergo by making it faster, less invasive, and more affordable. Pathological changes can begin 15-20 years before cognitive symptoms are clearly noticeable to others Jack et al. (2010), so a quicker and more sensitive detection mechanism could enable earlier intervention. This may potentially slow disease progression and delay costly care, projected to reach $7 trillion in the U.S. by 2050.222https://www.prnewswire.com/news-releases/new-alzheimers-association-report-reveals-sharp-increases-in-alzheimers-prevalence-deaths-and-cost-of-care-300616178.html
The second motivation to pursue dementia detection is a data-driven one. NLP research is inherently data-driven, and in this domain, data is often structured—spanning medical, audio, or textual modalities—annotated with categorical labels such as ’healthy’ or ’dementia.’ This structure perfectly suits classification algorithms, making dementia detection a well-defined challenge for researchers.
NLP-based dementia detection is a research domain spanning over a decade, primarily relying on well-known datasets from the DementiaBank cohort Lanzi et al. (2023). Most used is the Pitt corpus Becker et al. (1994), offering transcribed recordings of cognitive assessments such as picture descriptions. The popular ADReSS and ADReSSo challenge datasets (Luz et al., 2021a, b, respectively) are derivatives of the Pitt corpus, offering refined transcripts and a demographically balanced sample. The CCC dataset Pope and Davis (2011) contains transcribed conversations on memories, daily life, and health. Most studies on detection utilize these transcribed conversations (or alternatively, use self-collected clinical notes), annotated with Dementia or Healthy Control labels.333Some datasets provide more granular annotations such as Dementia, Mild Cognitive Impairment, Alzheimer’s, etc.
Across the 116 detection papers we reviewed, several algorithmic approaches emerged. Please note that we present representative citations here; for the complete list of references, see Appendix B. Before 2017, many studies relied on classic machine learning methods such as SVM and Random Forests using straightforward features like N-grams (Jarrold et al., 2014; Fraser et al., 2016a; Zhou et al., 2016; SANTOS et al., 2017), achieving accuracy rates around 85%. From 2018 onwards, there was a surge in transformer-based classifiers, particularly applied to the ADReSS and ADReSSo challenges published around that time (Pappagari et al., 2020; Edwards et al., 2020; Haulcy and Glass, 2021; Balagopalan et al., 2021). These methods pushed accuracy over 90%, representing the current state-of-the-art on the Pitt corpus and its subsets. Recently, large language models (LLMs) have been utilized for detection via prompt design, embedding extraction, and feature design (Agbavor and Liang, 2022; Wang et al., 2023a; Runde et al., 2024; BT and Chen, 2024; Bang et al., 2024).
Given the extensive body of works and impressive classification results, one might assume that dementia detection is a solved task. However, (a) these results are mostly based on the Pitt corpus or its subsets—small, homogeneous datasets with very specific characteristics (structured conversation, English only, etc.); (b) only 24% of these works report statistical significance or assess the robustness of their findings; and (c) despite strong results, no NLP-based classification tool, to our knowledge, has advanced to a real-world application. We explore these gaps further in Section 4.
Linguistic Bio-Markers Extraction
Some researchers use NLP not for detection, but for various types of linguistic exploration. One motivation is to confirm, on a large scale, the existing knowledge we have about the language of cognitively impaired individuals (as reviewed by Clark et al., 2014; Fraser et al., 2016a; Voleti et al., 2019, and others). For instance, studies such as Orimaye et al. (2014), Rosas et al. (2019) and Ilias and Askounis (2022a) used NLP methods to solidify that linguistic behaviors associated with cognitive decline—such as repetitions, revisions, and simpler language—are much more common in cognitively impaired individuals than in healthy controls.
Other studies introduce novel linguistic metrics and markers to assess a speaker’s cognitive state. Examples include complexity scores from parse trees (Roark et al., 2007), idea density, which measures the efficiency of idea conveyance (Sirts et al., 2017), and the use of meta-semantic terms in picture descriptions, such as words implying emotion, emphasis, or opinion (Choi et al., 2019). The amount of disfluencies—including silent pauses, reformulations, and context switches (Adhikari et al., 2021; Farzana et al., 2022; Williams et al., 2023)—have also proven to be strong linguistic markers, even enabling the longitudinal tracking of disease progression (Martinc et al., 2021; Robin et al., 2023). These findings may influence NLP practices as they show that certain preprocessing steps, like removing repetitions and stop words, could erase valuable dementia-related signals. They may also impact the medical community by suggesting new metrics, like idea density, for novel cognitive assessments.
Notably, 70% of works in this task family report statistical significance—the highest ratio among all task families. This likely stems from the motivation to ensure that existing knowledge, or any new linguistic insights, are robust.
Caregiver Support
With the growing number of dementia patients, the demand for caregivers rises. In the US alone, an estimated 11 million Americans serve as caregivers for those with dementia, contributing billions of hours of care each year.444https://www.alzint.org/about/dementia-facts-figures/ Research indicates that nearly half of these caregivers experience depression, and are at a higher risk of chronic health conditions.555https://www.caregiver.org/news/caregiver-isolation-and-loneliness/ NLP methods can help support these devoted caregivers by detecting emotional distress, providing answers to their concerns and offering companionship.
NLP for caregiver support, comprising less than 6% of our reviewed works, emerged after the COVID-19 outbreak, focusing on the emotional well-being of personal caregivers—i.e., family and friends- primarily through social media posts (Monfared et al., 2021). For example, Sunmoo et al. (2022, 2023) used graph-based topic modeling and sentiment analysis to show a shift in tweets posted throughout the pandemic—from practical care topics to emotional distress (e.g., depression, helplessness, elder abuse) and coping strategies (e.g., therapeutic reading). Posts beyond dementia-specific communities (Ni et al., 2022) and across social media profiles (Klein et al., 2022) reveal that caregivers express more than emotional distress, asking also for financial aid and legal advice.
Focusing on professional caregivers, Zhu et al. (2022b) analyzed clinical notes by nurses in aged care facilities. Disturbingly, they unveil dozens of distinct aggressive behaviors—including pushing, shouting, and using profane language—involving over 50% of dementia patients studied. Works like these highlight the potential of NLP in detecting when to assist caregivers, whether through advice, emotional support, or enhanced workforce training.
A growing body of research is exploring the potential of LLMs in providing caregivers with answers to dementia-related inquiries. Current studies show that models like ChatGPT offer relevant and factually correct advice and recommendations (Hristidis et al., 2023; Dosso et al., 2024), but often lack the depth of information available through Google Search or expert sources. Furthermore, LLMs fall short in offering the emotional support caregivers need when seeking to manage a loved one’s memory loss, confusion, or aggression (Aguirre et al., 2024). To bridge these gaps, Zaman et al. (2023) and Parmanto et al. (2024) explore the novel concept of LLMs fine-tuned specifically to address caregivers’ delicate emotional needs. These studies take a step toward AI-based caregiver companions, a field that is likely to gain traction with the rising popularity of LLMs.
Patient Assistance
Individuals with dementia can live with the disease for up to 20 years,666https://www.alzheimers.org.uk/about-dementia/symptoms-and-diagnosis/how-dementia-progresses/later-stages-dementia and NLP solutions can significantly improve their daily lives. NLP can also help researchers decode the complex mental landscapes of patients, where agitation, depression, and cognitive decline intersect with daily experiences. Research indicates that depression affects 30-50% of dementia patients Lyketsos and Lee (2003), with 15% experiencing suicidal thoughts Naismith et al. (2022). This highlights the importance of works such as Fraser et al. (2016b) and Ehghaghi et al. (2022), who aim to detect depression among dementia patients. These studies reveal that relying solely on transcriptions of cognitive assessments makes it extremely difficult to detect depression among dementia patients, and that incorporating multi-modal data, particularly audio, enhances the results.
Apart from assessing the mental well-being of patients, NLP can also aid in text simplification, making online information more accessible to those who often find it unreadable Espinosa-Zaragoza et al. (2023). Engineer et al. (2023) examined dementia-related texts online, revealing that the average readability level corresponds to 12 years of education—a level that many communities at higher risk for dementia may not have, given the research linking lower education levels to increased dementia susceptibility.777https://www.ncbi.nlm.nih.gov/pmc/articles/PMC6937498/ Additionally, the study reveals that the texts have a consistently negative tone, potentially affecting the mental state of patients who are already prone to depression. NLP tools are natural candidates to assist writers and content creators in producing more readable and supportive content.
Researchers are also working to help patients sustain meaningful conversations and prevent communication breakdowns, shown to occur in over a third of interactions involving individuals with dementia, as shown by Chinaei et al. (2017). Both they and Green et al. (2012) have developed systems based on transcribed interviews to detect confusion and disfluent speech patterns, applying real-time repair strategies to reduce misunderstandings. These systems enhance the conversational experiences of dementia patients, improving their quality of life and reducing the burden on caregivers.
A growing body of research is exploring the potential of LLM-powered chatbots to improve social engagement, promote cognitive stimulation, and alleviate loneliness. Xygkou et al. (2024) and Qi (2023) explore GPT-based agents for reminiscence therapy (engaging in conversations about personal photos) as well as for general social interaction and accompaniment. Both studies found that participants valued the novelty and cognitive stimulation, reporting reduced loneliness and increased social support. However, participants noted limitations in handling emotionally sensitive conversations.
LLMs for patients appear to be the next frontier in this task family, and many milestones have yet to be achieved—improved emotional capabilities, larger-scale experiments, and the development of evaluation benchmarks Gholizadeh et al. (2023). Additionally, as with all AI tools, building trust with patients and caregivers is essential. Gilman et al. (2024) proposes specialized GPT training for dementia patients as a way to help establish this trust. Treder et al. (2024) interviewed patients and caregivers about their views on such technologies, revealing concerns about bias, data privacy, and emotional intelligence, alongside a heartwarming, cautious optimism. As one patient experimenting with a chatbot noted: "I could talk to the robot longer than I could talk to a human… she didn’t tell me if I’ve repeated myself… she [the bot] didn’t think I was boring". Xygkou et al. (2024)
Summary
Figure 3 illustrates the contribution of various scientific communities to the research areas discussed above, including literature reviews and dataset introductions. While all communities participate in research on dementia detection as well as publish new datasets, the field of caregiver support has not yet been explored by the NLP and Speech communities. Additionally, we found no literature review similar to ours in NLP venues, highlighting a significant gap. We explore these and other gaps in the next section, pointing out opportunities for deeper, more rigorous research and innovative, exciting directions.
4 Gaps and Future Research
4.1 Data Variety
Classic
We begin by addressing the core of NLP research: the data. As noted in Section 3, the most popular datasets in the field are the Pitt corpus and its derivatives (ADReSS and ADReSSo), along with the CCC dataset (see additional details in Table 1). They are foundational and have significantly shaped the current state of NLP for dementia. However, like all data sources, they have limitations, including size constraints and demographic biases. For instance, the Pitt corpus contains more women than men and an imbalanced ratio of 35% healthy, 48% dementia, and 17% likely dementia participants. The ADReSS dataset was created as a balanced subset of Pitt, offering a smaller but evenly split sample by age, gender, and diagnosis Ševčík and Rusko (2022). Additionally, these datasets capture only the language of individuals with the access, ability, and willingness to participate in clinical trials. As a result, communities with less access to such trials, limited education, or intellectual and physical disabilities—factors known to impact cognitive assessments—are not represented Bruhn and Dammeyer (2018).
60% of dementia patients are non-English speakers from low- and middle-income countries,888https://www.alzint.org/about/dementia-facts-figures/ making it crucial to gather data from these communities. Some studies try to bridge this data gap using translation methods and transfer learning for dementia detection in resource-limited languages (Drame et al., 2012; Fraser et al., 2019; Guo et al., 2020; Pérez-Toro et al., 2023, 2022; Kabir et al., 2023; Meng et al., 2023; Melistas et al., 2023). Nowenstein et al. (2024), who applied translation for dementia detection in Icelandic, caution that despite their promise, such methods need extensive validation due to varying grammatical complexities and dementia-related nuances across languages.
Another concern regarding these classic datasets is their timeliness. For example, in the Pitt Corpus, published in 1994, the median participant age was 67, i.e., these individuals would be approximately 97 years old today. There is a significant difference between 67-year-olds now versus 30 years ago, including changes in technological proficiency and language use. This raises questions about whether models trained on these datasets can still provide accurate early diagnosis for today’s elderly population and whether the datasets could potentially be reproduced with updated protocols.
Another limitation of these datasets, and any dataset composed of transcribed audio, is their sensitivity to variations in automatic speech recognition (ASR) systems (Al Hanai et al., 2018; Balagopalan et al., 2019; Gómez-Zaragozá et al., 2023; Heitz et al., 2024). This issue is particularly critical for dementia-related data, where important cognitive indicators such as repetitions and silent pauses can be lost with certain ASR techniques.
Lastly, while these classic datasets significantly support research on tasks like detection and linguistic marker extraction, they offer no insight into caregivers, for example. They focus exclusively on patient data, overlooking other dementia stakeholders (Tahami Monfared et al., 2022), including the person conducting the cognitive assessment.
Contemporary
Let’s explore data sources beyond transcribed clinical trials. Publicly available data, like social media, has proven valuable for capturing the broad spectrum of communication across all dementia stakeholders (as discussed in Section 3). Additionally, some works creatively utilize non-dementia datasets like IMDB, leveraging its size and multilingual properties for translation methods and embedding extraction (Mirheidari et al., 2018; Li et al., 2019).
| Name | Year | Type | Population | Modalities | Size | |||||
| Pitt | 1994 | Cognitive Assessments | Patients | Transcribed speech | 196 Dem vs. 98 Ctrl | |||||
| ADReSS | 2020 | Cognitive Assessments | Patients | Transcribed speech | 78 Dem vs. 78 Ctrl | |||||
| ADReSSo | 2021 | Cognitive Assessments | Patients | Transcribed speech | 121 Dem vs. 116 Ctrl | |||||
| Classic | CCC | 2011 |
Patients | Transcribed speech | 125 Dem vs.125 Ctrl | ||||
| CareD | 2023 | Social media posts | Caregivers | Caregivers written text | 1005 Posts | |||||
| LoSST-AD | 2024 | Public interviews | Patients (Celebrities) | Transcribed Interviews | 10 Dem vs. 10 Ctrl | |||||
| Contemporary | Li et-al | 2023 |
Clinicians / Patients |
30k sentences |
An additional direction is freely available data of public figures diagnosed with dementia. Petti et al. (2023a, b) and Asllani and Mullen (2023) examine personal writings and spontaneous speech from well-known individuals with dementia, while Petti and Korhonen (2024) curated a corpus of a similar nature (see Appendix C). This data source is unique in spanning celebrities of different cultures and origins, different types of data sources (interviews, written texts, etc.), and different modalities (video, audio, etc.). Additionally, it allows for tracking these public figures throughout their disease journey, exploring their progression in an individualized way. Berisha et al. (2015) and Wang et al. (2020) emphasize this in their detailed case studies on the linguistic decline of President Ronald Reagan throughout his political career.
LLMs offer another promising approach to data collection, annotation, and augmentation, with proven success in enhancing clinical notes (Liu et al., 2023a; Koga et al., 2024; Latif and Kim, 2024). In a unique study, Li et al. (2023) showed that LLMs can label clinical records, generate clinical data, and spontaneous speech, improving detection performance beyond that of purely human-created datasets. The datasets they present are, to the best of our knowledge, the only dementia-related synthetic datasets currently available.
Future
Historically, ’NLP for dementia’ was synonymous with dementia detection, and annotated clinical records have proven well-suited for detection classifiers. However, to advance dementia research in all tasks and frontiers, NLP requires data that is abundant, diverse, and reflective of the disease’s complexity. Table 1 compares the traditional clinical datasets with the more contemporary ones we discussed, representing a small part of the 17 datasets we identified in our review (Table 2 in Appendix C). We hope these resources inspire researchers to pursue data from creative avenues, covering diverse dementia stakeholders, multiple languages, genders, cultures, and educational backgrounds, designing NLP solutions that are accessible to all.
4.2 Personalized Approaches
Every dementia patient experiences the disease’s progression uniquely, shaped by their individual characteristics and distinct personal language model. As each generation produces more accessible data, a 60-year-old today has likely generated significantly more text than a 60-year-old two decades ago. Given this wealth of textual data, can NLP facilitate personalized approaches for managing, tracking, or even diagnosing dementia by building individualized language models that monitor cognition on a personal level?
NLP-based digital twins for dementia patients represent a promising frontier in healthcare. These virtual models replicate a patient’s cognitive and behavioral patterns through their language, enabling personalized care by monitoring cognitive decline and adjusting treatment plans based on simulated behaviors Andargoli et al. (2024). A notable example is Hong et al. (2024), who developed digital twins by fine-tuning an LLM on longitudinal data from the I-CONECT clinical trial,999https://www.i-conect.org/ using these twins to evaluate a patient support chatbot.
An intriguing, yet unexplored, research direction involves creating personalized LLM companions modeled after loved ones, as cognitive stimuli from familiar individuals can positively impact dementia patients Woods et al. (2012). An LLM trained on family-curated data could engage patients with their personal history, family stories, and interests—such as favorite bands or recipes—potentially offering a non-clinical treatment. Given the vast amount of personal data generated today, this approach seems both feasible and effective. With these research directions, NLP has the potential to shape the future of personalized dementia treatments, both medical and alternative.
4.3 Trust and Scientific Rigor
A key contribution of our work lies in our ability to draw comparisons between the medical and technological communities, differing in various ways—one of which is scientific modeling. Medical researchers typically follow a hypothesis-driven process, designing studies and collecting data to test specific assumptions. In contrast, data scientists often retrospectively explore pre-collected data, searching for patterns without necessarily defining a specific hypothesis. This raises concerns for medical researchers: is the retrospective data relevant to future contexts (e.g., will a model trained on the Pitt corpus from 1994 capture the same signals in today’s dementia patients)? Can the model generalize well? Do the findings highlight correlation rather than causation? Indeed, the medical community, with its practical concerns, might be apprehensive about adopting solutions that, in their view, extract correlations from retrospectively collected, often small datasets, as this approach contrasts significantly with their scientific view of research. Their apprehension might be exacerbated by any lack of reporting on statistical significance or evaluation of robustness.
To assess this, we annotated papers based on whether they report p-values or statistical significance tests, using it as a proxy for scientific rigor in NLP for dementia research.101010For literature reviews, statistical significance was marked as not relevant. Among all reviewed works, only 30% report statistical significance. While nearly half of NLP and medical papers include it, around 86% of works published in Speech venues —covering a significant portion of detection studies, particularly those using the popular ADReSS and ADReSSo datasets—do not.
Another crucial aspect of scientific modeling, alongside statistical significance, is robustness. One work addressing this is Novikova (2021), who examines the robustness of BERT-based dementia detection classifiers to data perturbations. Another example is Favaro et al. (2024), who investigates the consistency of linguistic features across multiple annotated corpora, identifying features that reliably differentiate between dementia and control groups while correlating with clinical scores.
A third aspect of scientific rigor is the adoption of explainability mechanisms in NLP-based solutions. This is particularly crucial for dementia detection algorithms that may eventually serve as real-world decision-support systems for doctors. Although some detection works prioritize explainable mechanisms (Ilias and Askounis, 2022a), such works remain relatively rare. The medical community is generally open to AI-based decision-support systems but demands transparency, understandably mistrusting "black-box" models (Adadi and Berrada, 2020). In fact, while the NLP community focuses on hard-to-decipher deep learning models—due to their higher detection accuracy—the medical community may be more inclined to adopt traditional machine learning classifiers given their transparency and interpretability.
Ensuring scientific rigor is equally crucial—and challenging—for other task families. Works on caregiver support and patient assistance often lack a gold standard or clear objective, unlike the well-defined task of dementia detection. For example, evaluating LLMs that answer medical questions or simplify texts for dementia patients presents difficulties in terms of assessment and statistical rigor. Yet, understanding model outputs, assessing robustness, and clarifying motivations remain essential. Hence, developing benchmarks to evaluate these tools is a vital research direction.
Without transparent reporting of significance results, a focus on robustness and explainability, and clear evaluation benchmarks, NLP for dementia is at risk of remaining a tool limited to academic discussions, making its adoption into the medical community and other real-world applications significantly more challenging.
4.4 Application State-of-Mind
Alongside adopting rigorous methodologies and establishing trust, combating dementia in real-world settings requires NLP practitioners to embrace an application-focused mindset—’how can my work be applied in practice?’ A prominent example arises from our review of over 110 studies on NLP-driven dementia detection. Despite decades of promising results, we found no evidence in our research that any of these approaches have been integrated into real-world detection processes.111111We base this on publicly available data but acknowledge that efforts may exist beyond public knowledge
This raises the question: what is preventing the application of these promising NLP techniques in clinical settings? We hypothesize that trust issues (as discussed in the previous subsection) play a significant role, as it is simply too risky to adopt such tools into medical processes without rigorous testing across multiple datasets, contexts, and real-world scenarios. We believe that if researchers intentionally design their solutions with real-world applications in mind, outcomes may be different.
Applicative thinking goes beyond dementia detection—NLP tools for various tasks can be applied in practice with the right mindset, cross-community collaboration, and rigorous testing. For example, Topaz et al. (2020) developed a system to help home healthcare providers identify and cluster neuropsychiatric symptoms (e.g., anxiety, delusions), flagging dementia-related signs even in undiagnosed individuals. This system supports information transfer across healthcare settings and aids both professional and personal caregivers.
Demand for dementia-related applications is evident from both business and technology perspectives. The dementia management and treatment market is expected to exceed a value of $36 billion USD by 2030,121212https://www.sphericalinsights.com/reports/dementia-drugs-market and independent technological solutions are quickly emerging. For instance, the ChatGPT ’app store’ now features over 30 dementia-related chatbots for cognitive stimulation or virtual companionship.131313Searched on https://chat.openai.com/, August 2024 People are taking initiative, leveraging LLMs to create these tools, and scientists can follow suit, using this framework to develop LLM-based playgrounds for large-scale applicative experiments. The need for practical NLP tools is undeniable, and we encourage the community to focus on solutions that are not only innovative but also truly applicable.
4.5 NLP Insights for Medicine
This work highlights the efforts of diverse scientific communities united by the common goal of advancing dementia research, and cross-disciplinary knowledge-sharing is imperative for success. For example, medical experts can collaborate with NLP practitioners when building datasets, while NLP researchers should work with medical professionals to develop robust and explainable solutions.
Knowledge transfer between the medical and technological communities should flow both ways. Currently, it is primarily the technological community (especially NLP) building on medical, linguistic, and cognitive knowledge in dementia research—for instance, using established linguistic markers and datasets originally collected through medical experiments like cognitive assessments. However, we argue that the other direction—where the medical community draws insights from NLP research—holds transformative potential.
A specific example is the development of artificially degraded language models, which explore ’healthy’ vs. ’cognitively impaired’ models that replicate the linguistic behavior of dementia patients. Cohen and Pakhomov (2020) achieved this by fine-tuning an LSTM on the DementiaBank cohort, while Li et al. (2022) degraded different layers of GPT-2. Both studies impressively demonstrated dementia-like linguistic patterns in their models. Additionally, these works introduced perplexity as a useful metric for measuring cognitive decline, showing that degraded models exhibited significantly lower perplexity when exposed to dementia-related linguistic anomalies compared to ’healthy’ models. We believe that this research area can help us better understand the biological mechanisms of dementia and its pathologies. For example, an interesting direction would be to artificially degrade an LLM by mimicking protein deposits on neurons, potentially offering insights into the mechanisms of Alzheimer’s disease.
Loconte et al. (2023), on the other hand, question whether LLMs can truly enhance our understanding of the brain in the context of dementia. Testing ChatGPT with neuropsychological assessments for prefrontal functions, they find inconsistent results, showing these models still fall short of mimicking human cognition. This research sheds light on both cognitive mechanisms and how well neural networks model our brain’s complexity.
It’s not only the medical or biological community that could learn from NLP research. For example, the work of Li et al. (2024), which examines the parallels between the language of children and people with dementia, may advance research in cognitive sciences and linguistics. Using their artificially degraded language model mentioned above, they show that it is significantly less perplexed than GPT-2 when analyzing picture descriptions by young children and dementia patients. Looking at children above the age of 8, whose language aligns more closely with healthy adults, GPT-2 becomes less perplexed. This study highlights the potential for designing dementia-related biomarkers or novel cognitive assessments by learning from the language patterns of children.
We believe NLP has great potential for developing new cognitive assessment tests. Several studies suggest focusing on the dynamics between the interviewer and patient, rather than just the task at hand. They demonstrate that rare dialogue acts, such as clarification exchanges, and the interviewer’s role in the conversation can reveal cognitive signals that can potentially inform the creation of novel assessments (Farzana et al., 2020; Nasreen et al., 2021a; Farzana and Parde, 2022; Dawalatabad et al., 2022). Another example is Reeves et al. (2020), who introduced a novel NLP-empowered cognitive test based on narrative descriptions after watching video clips.
Finally, NLP research may even prompt medical professionals to reconsider existing hypotheses about dementia risk factors. Zhou et al. (2019) analyzed a large cohort of text-rich health records and found a contradiction with contemporary studies that emphasize cardiovascular risks, such as high-fat, high-calorie, and high-carbohydrate diets, as causes of dementia. Surprisingly, these factors were rarely mentioned in patients’ clinical records. Instead, they found significant associations with deficiencies in potassium, calcium, and other nutrients, mentioned in over 25% of dementia patients’ records. However, actual laboratory results for nutrient levels were rarely documented. This study highlights NLP’s potential to challenge existing hypotheses, and expose gaps in medical procedures.
5 Conclusions
Our review of 205 papers on NLP for dementia, spanning multiple scientific communities, reveals a richness of tasks, datasets, gaps, and creative opportunities. We show that NLP for dementia extends far beyond detection, with the potential to transform the lives of caregivers and patients, while shaping medical research on the disease. We call on researchers to collaborate closely and share knowledge through rigorous research, so that together we can make a meaningful impact on people’s lives and combat this disease.
Appendix A Cohort Construction
A.1 Search Methodology
We searched in six data sources: ACL Anthology, PubMed, DBLP, IEEE Xplore, Springer, and Wiley. The first database search was conducted on April 10, 2024, and the last one on July 19, 2024. Our research strategy focused on papers addressing dementia, Alzheimer’s, and Mild Cognitive Impairment (MCI) using NLP, including but not limited to large language models (LLMs). We searched for papers including the keywords ("NLP" OR "Natural Language Processing" OR "LLM" OR "Language Model" OR "ChatGPT") AND ("Dementia" OR "Alzheimer" OR "MCI" OR "Cognitive Impairment") in their title and/or abstract. Although we don’t include keywords for specific dementia pathologies (e.g., ’Lewy Body’), we believe that using the general term ’dementia’ in the title and/or abstract will capture papers on various types.
A.2 Paper inclusion and Exclusion Criteria
We screened the extracted articles against the following criteria:
a)
Works applying NLP technologies, algorithms, and methodologies to textual data (including speech transcriptions) were included.
b)
Multi-modal approaches, as long as they include some form of textual analysis, were included.
c)
Papers focusing solely on audio, visual, or other non-textual data were excluded.
d)
Only fully accessible, English-language papers were included.
e)
Only peer-reviewed papers were considered.
A.3 Dataset inclusion and Exclusion Criteria
We screened the featured datasets against the following criteria:
a)
Only datasets mentioned in the reviewed works were included.
b)
Textual data must be fully or partially in English.
c)
Sufficient information about the dataset must be available online or through private requests from researchers or institutions.
d)
Data does not need to be publicly accessible to be included in our review.
e)
As these datasets are part of the peer-reviewed literature we selected, we assume they have also undergone peer review.
A.4 Paper Annotation
For each paper, we extracted the title, authors, year of publication, and venue. We then manually annotated the following:
a)
Task Family: Dementia detection, linguistic bio-marker extraction, caregiver support, patient assistance, literature review, or dataset introduction. While some papers may address multiple categories, we selected the most prominent motivation and novelty as described by the authors.
b)
Venue type: NLP (e.g., ACL), Medical (e.g., NIH), Speech (e.g., InterSPEECH), or Other Technological (e.g., Frontiers in CS)
c)
Datasets used (if applicable).
d)
Technologies applied (if applicable), such as SVM or BERT.
e)
Statistical significance reported: yes or no, depending on whether any statistical tests were conducted on data or algorithmic results.
Appendix B Extended Mention of Works
In Section 3, we cited a representative sample of dementia detection works we reviewed. Below, we cite the remainder of the papers according to internal categorization used for our internal analysis.
Detection works using classic ML classifiers on a wide range of features, from straightforward N-grams and information content units extracted from picture descriptions to dementia patients’ dental records and motor signs: Prud’hommeaux and Roark (2011), Fraser et al. (2013), Jarrold et al. (2014), Prud’hommeaux and Roark (2015), Fraser et al. (2016a), Yancheva and Rudzicz (2016), Zhou et al. (2016), Weissenbacher et al. (2016), Mirheidari et al. (2016), SANTOS et al. (2017), Wankerl et al. (2017), Pou-Prom and Rudzicz (2018), Al-Harrasi et al. (2021), Eyre et al. (2020), Chen et al. (2020), Searle et al. (2020), Martinc and Pollak (2020), Hane et al. (2020), Nasreen et al. (2021b), Gónzalez Atienza et al. (2021), Soni et al. (2021), Penfold et al. (2022), Ablimit et al. (2022b), Amini et al. (2023), Taghibeyglou and Rudzicz (2023), Patel and Wu (2024).
Works using classic neural models such as LSTMs and transformer-based models like BERT: Karlekar et al. (2018), Pompili et al. (2018), Pou-Prom and Rudzicz (2018), Di Palo and Parde (2019), Chen et al. (2019), Edwards et al. (2020), Syed et al. (2020), Koo et al. (2020), Sarawgi et al. (2020), Pompili et al. (2020), Cummins et al. (2020), Balagopalan et al. (2020), Yuan et al. (2020), Pappagari et al. (2020), Haulcy and Glass (2021), Balagopalan et al. (2021), Wang et al. (2021), Qiao et al. (2021), Campbell et al. (2021), Rohanian et al. (2021a), Syed et al. (2021), Pappagari et al. (2021), Pérez-Toro et al. (2021), Mirheidari et al. (2021), Liu et al. (2021), Zhu et al. (2021), Saltz et al. (2021), Khan et al. (2022), Deng et al. (2022), Horigome et al. (2022), Zheng et al. (2022), Liu et al. (2022b), Liu et al. (2022a), Wang et al. (2022), Matošević and Jović (2022), Amini et al. (2023), Abdelhalim et al. (2023), Bouazizi et al. (2023), Gkoumas et al. (2023), Cai et al. (2023), Wang et al. (2023b), Shakeri et al. (2024), Dong et al. (2024), Kim et al. (2024).
Works classifying between more granular categories such as ’MCI’, ’Alzheimer’s’, and ’Probable Alzheimer’s’: Mao et al. (2023), Liu et al. (2023b), Amini et al. (2024), Laurentiev et al. (2024). Works focusing on text-based MMSE score regression for dementia detection: Yancheva et al. (2015), Pou-Prom and Rudzicz (2018), Rohanian et al. (2021b).
Works presenting multi-modal approaches (with audio / video) as well as augmentation methods using synthetic, noisy, cross-domain, or semantically and structurally different datasets: Che et al. (2017), Gligic et al. (2020), Gupta et al. (2020), Amin-Nejad et al. (2020), Guo et al. (2021), Zhu et al. (2021), Zhu et al. (2022a), Ablimit et al. (2022a), Mirheidari et al. (2022), Hlédiková et al. (2022), Ilias and Askounis (2022b), Maclagan et al. (2023), Chen et al. (2023b), Farzana and Parde (2023), Cui et al. (2023), Chen et al. (2023a), Cai et al. (2023), Lin and Washington (2024), Fard et al. (2024). Works leveraging LLMs for data augmentation: Bullard et al. (2016), Liu et al. (2019), Chintagunta et al. (2021), Liu et al. (2023a), Cai et al. (2023), Duan et al. (2023), Koga et al. (2024), Latif and Kim (2024).
Appendix C Full Dataset Table
Table 2 (presented in a separate page due to margin requirements of the TACL format) showcases the dementia-related datasets encountered throughout our review. Datasets such as the Reagan Library and IMDB, though creatively used in some works, are excluded as they are not dementia-specific. Private health records or unreleased clinical notes are excluded as well.
| General | Data Source | Modalities | Participants | ||||||||||||||
| Dataset | Longit- | Clinical | Social | Non-clinic. | Synt. | Trans. | Written | Audio | Video | Medical | Physical | Patients | Other | ||||
| udinal | setting | media | interviews | data | speech | text | imaging | markers | |||||||||
x | x | - | - | - | x | - | x | - | - | - |
- | x | - | - | - | x | - | x | - | - | - |
- | x | - | - | - | - | - | x | - | - | - |
x | - | - | x | - | x | - | x | x | - | - |
x | x | - | - | - | x | - | x | - | - | - |
- | x | - | - | - | - | - | x | - | - | - |
| Farmington Heart Study | x | x | - | - | - | - | x | x | - | x | x |
- | - | x | - | - | - | x | x | x | x | x | - |
x | x | - | - | - | - | x | - | - | x | x |
x | x | - | x | - | x | x | - | - | - | x |
x | x | - | - | - | - | x | - | - | x | x | >4800 | - | ||||
- | x | - | x | - | x | - | x | x | x | x |
| I-CONECT | x | - | - | x | - | - | - | - | x | - | - | 320* | - | ||||
x | - | - | x | - | x | - | - | - | - | - |
x | x | - | - | - | x | - | x | - | - | x |
| Li et al. (2023) | - | x | - | - | x | - | x | - | - | - | - | - |
| Gkoumas et al. (2024) | x | - | - | x | - | x | x | x | - | - | - |
** Estimated number of participants in the I-CONECT clinical trial are based on publicly available data.